OpenStudy (anonymous):

Naturally occurring copper has two isotopes with relative atomic mass 62.930 and 64.928 a.m.u. The average relative atomic mass of copper is 63.546. What are the relative abundances of the two isotopes in naturally occurring copper?

OpenStudy (ookawaiioo):

I will give you an equation........

OpenStudy (ookawaiioo):

63.546 = (62.930(X)+64.928(Y))/2 X+Y= 1 X and Y are less than 1, but add up to one. Once you find the X and Y, multiple each by 100 to get a percent. Goodluck.

OpenStudy (anonymous):

62.930X + 64.928Y = 63.546 62.930X + 64.546(1 - X) = 63.546 X + Y = 1 62.930X + 64.928 -64.928X = 63.546 62.930X - 64.928X = 63.546 - 64.928 like terms together -1.998X = -1.382 \[X = \frac{ -1.382 }{ -1.998 }\] X = 0.691691691 X = 0.69 x 100 = 69% Y =1 - 0.69 = 0.31 = 0.31 x 100 = 31%